OpenStudy (anonymous):

Prove that sqrt2is irrational.

mathslover (mathslover):

ok so what is an irrational number?

Parth (parthkohli):

@higgs Proof by contradiction!

mathslover (mathslover):

Irrational number is a number which can not be represented in the form of p/q where p and q are integers and q is not equal to zero

OpenStudy (anonymous):

Prove not explain

Parth (parthkohli):

First, assume that \(\sqrt2\) is rational. Therefore,\[ \sqrt{2} = {p \over q} \qquad[\gcd({p,q = 0)]}\]

mathslover (mathslover):

Let sqrt 2 be a rational number since sqrt 2 is taken as a rational number hence it must be represented in the form of p/q

Parth (parthkohli):

Square both sides.\[2 = {p^2 \over q^2}\]Multiply both sides by \(q^2\).\[ 2q^2 = p^2\]We already know that \(2q^2\) is an even number.

Parth (parthkohli):

We also know that \(p\) is even from the above right?

mathslover (mathslover):

oh ... wait : \[\large{\frac{a}{b}\times \frac{a}{b}=\frac{a^2}{b^2} = 2}\] \[\large{a^2=2b^2}\] since a^2 is even so a must be even let a = 2c or 4c^2 = 2b^2

Parth (parthkohli):

So, assume that \(p = 2k\).

Parth (parthkohli):

\[2 = {4k^2 \over q^2}\]

Parth (parthkohli):

\[2q^2 = 4k^2 \]

mathslover (mathslover):

or 2c^2=b^2 hence b must be even so a and b both are even .... but a/b was : reduced to lowest form hence sqrt{2} is not a rational number ... though it is an irrational number

mathslover (mathslover):

proved

Parth (parthkohli):

We already knew that \(p\) is even, but now that \(q\) is also even, we know that the gcd is 2!

mathslover (mathslover):

Note: there we assumed that HCF of a and b is 1

mathslover (mathslover):

got it @higgs ?

Parth (parthkohli):

But we first assumed that \(\gcd = 1\), but here's a contradiction!\[\gcd=2\Longleftrightarrow\gcd = 1\]

OpenStudy (anonymous):

I get it now thanks guys

mathslover (mathslover):

No problem

OpenStudy (zzr0ck3r):

lol, google!